OpenStudy (anonymous):

integrate xcos(x^2 + 1) using u substitution

jimthompson5910 (jim_thompson5910):

hint: u = x^2 + 1 du/dx = 2x du = 2x*dx du/2 = x*dx

OpenStudy (anonymous):

-sin(u)/2 , then sub, is that right?

jimthompson5910 (jim_thompson5910):

the integral of cos is not -sin

OpenStudy (anonymous):

oops, backwards, it's sin, right?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

sin(u)/2 and then yes, you replace u with x^2+1 don't forget about the +C

OpenStudy (anonymous):

always do, darn C

jimthompson5910 (jim_thompson5910):

yes since it's an indefinite integral

OpenStudy (anonymous):

ty jim